OpenStudy (anonymous):

Solve the equation \[\sin \theta = 2 \cos 2 \theta +1-------->0\le \theta \le 360\]

OpenStudy (anonymous):

sinx=2cos2x +1 sinx=2[cos ^2 x - sin ^2 x]+1 sinx=2[1-sin^2 x-sin^2 x]+1 sinx=2+1-4sin^2 x 4sin^2 x+sinx-3=0 4sin^2 x +4sinx-3sinx-3=0 4sinx(sinx+1)-3(sinx+1)=0 (4sinx-3)(sinx+1)=0 sinx=3/4 or sinx =-1 check it!!!

OpenStudy (anonymous):

oh sorry it was theta not x...

OpenStudy (anonymous):

very nice...

OpenStudy (anonymous):

I got to the bit where sinx=2+1-4sin^2 x, but i forgot how to carry on from there...

OpenStudy (anonymous):

just arrange it to the form of a quadratic equation in sinx

OpenStudy (anonymous):

did you get it @order???

OpenStudy (anonymous):

Yes, I'd just forgotten I could do that.